i am lost, is it possible to foil this, (3+h)^3......the third power is throwing me off here, any help would be greatly appreciated
Answer by nabla(475)

Foil the square first.
(3+h)^2=h^2+6h+9
Now, this is multiplied by (3+h)
(h^2+6h+9)(h+3)=h^3+3h^2+6h^2+18h+9h+27=h^3+9h^2+27h+27
The easiest way to remember how to do this is each term on the left gets multiplied by each term on the right...